If PQ is the double ordinate of the parabola y2=4ax then what is the locus of its point of trisection?

Grade:10

3 Answers

Arun
6 years ago
Dear Vijaya
 
The equation y^2 = 4ax can be parametrised as: 
x = at^2, y = 2at. 

If the extremities of the double ordinate are (at^2, 2at) and (at^2, - 2at), then its points of trisection have ordinates y1 and y2 where: 
y1 = - 2at + 4at / 3 = 2at / 3 
y2 = - 2at + 8at / 3 = - 2at / 3 

For both of these points: 
y^2 = 4a^2 t^2 / 9 
y^2 = (4a / 9)(at^2) 
y^2 = 4(a / 9)x. 

The locus is the parabola y^2 = 4bx where: 
b = a / 9.
hence locus -
y^2 = (4/9) ax
 
Regards
Arun(askIITians forum expert)
Harsh Vardhan
5 years ago
Let the extremities of double ordinate be (k,l) and (k,-l)
Let P(x1,y1) be any point on the double ordinate which divides the line in the ratio 1:2
 
P(x1,y1)=1(k)+2(k)/1+2,1(-1)+2(l)/1+2
=(K,l/3)
Here x1=k y1=l/3,l=3y1
(K,l)=(x1,3y1)
As (k,l) lies on the parabola
L^2=4ak
(3y1)^2=4ax1
9y1^2=4ax1
Hence the locus of the point of trisection is 
9y^2=4ax
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your problem.
 
The equation y^2 = 4ax can be parametrised as:
x = at^2, y = 2at.
 
If the extremities of the double ordinate are (at^2, 2at) and (at^2, - 2at), then its points of trisection have ordinates y1
and y2 where:
y1 = - 2at + 4at / 3 = 2at / 3
y2 = - 2at + 8at / 3 = - 2at / 3
 
For both of these points: 
y^2 = 4a^2 t^2 / 9 
y^2 = (4a / 9)(at^2) 
y^2 = 4(a / 9)x. 

The locus is the parabola y^2 = 4bx where: 
b = a / 9.
hence locus -
y^2 = (4/9) ax
 
Hope it helps.
Thanks and regards,
Kushagra